OpenStudy (anonymous):

Express in terms of sums and differences of logarithms, loga 3a/2b^4

OpenStudy (mertsj):

\[\log_{a}\frac{3a}{2b^4}=\log_{a}3a-\log_{a}2b^4=\log_{a}3+\log_{a}a-\log_{a}2-4\log_{a}b \]

OpenStudy (mertsj):

\[=\log_{a}3+1-\log_{a}2-4\log_{a}b \]